# On lower bounds for the distances between APN functions

### v1

@@ 5 Power functions
If $F(x)=x^{d}$ is a APN and $n$ is odd, we can determine the exact exclude multiplicities of $\mathcal{G}_{F}$ contained in $\left\{0\right\}\times\mathbb{F}_{2^{n}}^{\ast}$, but first, we describe the sum $\sum_{x\in\mathbb{F}_{2^{n}}}\gamma_{F}(x+a,x^{d}+b)$ in the following way to make our computation straightforward.
@@
In [27, Conjecture 21], Kaleyski conjectured that if $F$ is an APN power function, the two values $\sum_{y\in\mathbb{F}_{2^{n}}}\delta_{F}(y,F(y)+1)$ and $\sum_{y\in\mathbb{F}_{2^{n}}}\delta_{F}(y+1,F(y))$ are equal to $\sum_{y\in\mathbb{F}_{2^{n}}}\delta_{x^{3}}(y,y^{3}+1)$. By Proposition 3.3, this is equivalent to $\operatorname{mult}_{\mathcal{G}_{F}}(0,1)=\operatorname{mult}_{\mathcal{G}_{F}}(1,0)=\operatorname{mult}_{\mathcal{G}_{x^{3}}}(0,1)$. Therefore, Proposition 5.3 proves Kaleyski’s conjecture in the case of $n$ odd, but the case of $n$ even still remains open.
@@ 6 Lower bounds on distances from the APN inverse function
Kloosterman sums can also be used to describe the number of rational points on ordinary elliptic curves over $\mathbb{F}_{2^{n}}$. First, recall that an algebraic plane curve $\mathcal{C}$ over $\mathbb{F}_{2^{n}}$ is defined by the equation $p(x,y)=0$ for some irreducible polynomial $p(x,y)$ over $\mathbb{F}_{2^{n}}$. If $\widetilde{p}(x,y,z)$ is the associated homogenized polynomial to $p(x,y)$ we denote the solution set to $\widetilde{p}(x,y,z)=0$ as $\widetilde{C}$. We say that the number of rational points of $C$ is the size of $\widetilde{C}$. Also recall that a point $(x,y,z)\in\widetilde{C}$ is singular if all partial derivatives of $\widetilde{p}(x,y,z)$ vanish at $(x,y,z)$. If there are no singular points, we define the genus of $\mathcal{C}$ as $g=\frac{(d-1)(d-2)}{2}$, where $d$ is the degree of $\widetilde{p}(x,y,z)$. If the genus of $\mathcal{C}$ is $1$, then we say that $\mathcal{C}$ is elliptic. Moreover, we say that $\mathcal{C}$ is supersingular if the coefficient of xyzxyz is zero in $\widetilde{p}(x,y,z)$, otherwise we say that $\mathcal{C}$ is ordinary (cf. [32]).
@@
$\operatorname{mult}_{\mathcal{G}_{F}}(a,b)\geq\left\lceil\frac{2^{n}-5-\max_{a\in\mathbb{F}_{2^{n}}^{\ast}}|K_{n}(a)-1|}{6}\right\rceil$ (10)
@@
Let $\mathcal{E}=\{(x,y)\in\mathbb{F}_{2^{n}}^{2}:q(x,y)=0\}$, and let $G=\left\{(x,y)\in E:0\notin\left\{x,y,x+y+1\right\}\right\}$. Note that for all $(x,y)\in\mathbb{F}_{2^{n}}^{2}$ such that $x,y,x+y+1\neq 0$, we have $p(x,y)=0$ if and only if $q(x,y)=0.$ Hence $|D|=|G|$.
@@
Let $\widetilde{\mathcal{E}}=\{(x,y,z):\tilde{q}(x,y,z)=0\}$ denote the projective curve defined by $\tilde{q}$. By considering the partial derivatives of $\tilde{q}$, it is easy to see that $\mathcal{E}$ is smooth, and we know that $\mathcal{E}$ is elliptic because its genus is $g=\frac{(3-1)(3-2)}{2}=1$. Moreover, $\mathcal{E}$ is ordinary because xyzxyz has a nonzero coefficient in $\tilde{q}$. Note that the projective points of $\widetilde{\mathcal{E}}$ include three additional points, corresponding to solutions of the form $(x:y:0)$ that satisfy $\widetilde{q}(x,y,0)=bx^{2}y+bxy^{2}=bxy(x+y)=0$. There are three such cases: $x=0,y=0$ and $x+y=0,$ corresponding to the points $(1:0:0),(0:1:0),(1:1:0)\in\widetilde{E}$ respectively. So $\widetilde{\mathcal{E}}=\#\mathcal{E}+3$.
@@ 7 On the existence of linear structures of $\gamma_{F}$
In [13], there were two classes of APN functions that were shown that do not admit linear structures: the first if when $F$ is an APN power function, and the second is when $F$ is such that there exists some $x\in\mathbb{F}_{2}^{n}$ with the property that $\left\{(a,b)\in(\mathbb{F}_{2}^{n})^{2}:D_{a}D_{b}F(x)=w\right\}$ does not depend on $w\in\mathbb{F}_{2}^{n}\setminus\left\{0\right\}$.
We will prove that if $F$ is a plateaued 3-to-1 function ($n$ even), then $\gamma_{F}$ does not admit nontrivial linear structures, forming the third known class of APN functions such that $\gamma_{F}$ has no nontrivial linear structures. Let us rephrase the problem of $\gamma_{F}$ having linear structures in terms of exclude multiplicity when $F$ is a plateaued APN function.
Proposition 7.1.
@@
Proposition 7.1 shows that obtaining an upper bound on the exclude multiplicities of the graph of a plateaued APN function also poses as a difficult problem. However, recall that in Section 4.3, we determined the exact exclude multiplicities that occur when $F$ is a plateaued 3-to-1 function. The following result is then an immediate corollary.
@@
Assume $n\geq 4$ and is even. Let $F\colon\mathbb{F}_{2}^{n}\to\mathbb{F}_{2}^{n}$ be a plateaued 3-to-1 function. Then $\gamma_{F}$ has no nontrivial linear structures.
@@
By Theorem 4.20, any exclude point of $\mathcal{G}_{F}$ has multiplicity $\frac{2^{n}+(-2)^{\frac{n}{2}+1}-2}{6}$ or $\frac{2^{n}+(-2)^{\frac{n}{2}}-2}{6}$, and so any exclude point of $\mathcal{G}_{F}$ has multiplicity strictly less than $\frac{2^{n}-1}{3}$. The result then follows from Proposition 7.1. ∎
As shown above, there is an equivalence of $\gamma_{F}$ having a nontrivial linear structure and $\mathcal{G}_{F}$ having an exclude point of multiplicity $\frac{2^{n}-1}{3}$ when $F$ is plateaued APN. In the case of $F$ also being quadratic, we can then characterize the existence of a nontrivial linear structure of $\gamma_{F}$ in terms of the components of the ortho-derivative of $F$.
@@
Suppose $n\geq 4$ is even, and let $F\colon\mathbb{F}_{2}^{n}\to\mathbb{F}_{2}^{n}$ be a quadratic APN function. Then $\gamma_{F}$ has a nontrivial linear structure if and only if the ortho-derivative of $F$ has a component function of weight 0. Equivalently, $\gamma_{F}$ has a nontrivial linear structure if and only if the image of $\pi_{F}$ is contained in a linear hyperplane.
@@
Apply Proposition 7.1 to eq. 9. ∎
In particular, the class of quadratic APN functions whose ortho-derivatives have nonzero nonlinearity is a subclass of those APN functions such that $\gamma_{F}$ has no nontrivial linear structures. This immediately follows by applying the inequality $\operatorname{wt}(b\cdot F)\geq\operatorname{\mathcal{NL}}(\pi_{F})$ for $b\neq 0$ to eq. 9. In particular, we have the inequality
$\operatorname{mult}_{\mathcal{G}_{F}}(a,b)\leq\frac{2^{n}-1-\operatorname{\mathcal{NL}}(\pi_{F})}{3}.$ (13)
@@
Suppose $n\geq 4$ is even. Let $F\colon\mathbb{F}_{2}^{n}\to\mathbb{F}_{2}^{n}$ be a quadratic APN function with $F(0)=0$. Then for any $b\neq 0$, the weight of $D_{(0,b)}\gamma_{F}$ is divisible by $6\cdot 2^{n}$ and satisfies
@@
By Proposition 4.25, we know $\operatorname{mult}_{\mathcal{G}_{F}}(0,b)=2k+1$ for some non-negative integer $k\geq 0$. Therefore, by Proposition 4.17, we have $6\cdot 2^{n}(2k+1)=2^{n}(2^{n}-2)+2^{2n}-2\operatorname{wt}(D_{(0,b)}\gamma_{F})$. So $\operatorname{wt}(D_{(0,b)}\gamma_{F})=2^{n}(2^{n}-6k-4)$. Since $2^{n}-4$ is divisible by $6$ as $n$ is even, we have $\operatorname{wt}(D_{(0,b)}\gamma_{F})\equiv 0\mod 6\cdot 2^{n}$.
By Theorem 4.13, we have $\operatorname{mult}_{\mathcal{G}_{F}}(a,b)\geq 2^{\frac{n}{2}-1}-1$ for all $(a,b)\notin\mathcal{G}_{F}$ as $F$ is quadratic and therefore plateaued. Moreover, we have $\operatorname{mult}_{\mathcal{G}_{F}}(a,b)\leq\frac{2^{n}-1}{3}$ by [19, Corollary 3.4]. So, for $b\neq 0$, we have $2^{\frac{n}{2}-1}-1\leq\operatorname{mult}_{\mathcal{G}_{F}}(0,b)\leq\frac{2^{n}-1}{3}$, and the result follows. ∎
@@ 8 Open problems
For $n$ even, prove or disprove the existence of a plateaued (or quadratic) APN functions whose graph has an exclude point of multiplicity $\frac{2^{n}-1}{3}$ (the maximum that is theoretically possible). Recall that Proposition 7.1 characterizes this problem in terms of the existence of a nontrivial linear structure of $\gamma_{F}$.
@@
Open Problem 8.5.
For $n$ even, find an example of a non-quadratic APN function $F\colon\mathbb{F}_{2}^{n}\to\mathbb{F}_{2}^{n}$ whose graph has an exclude point of even multiplicity, or prove such a function does not exist. Note that if such a function does not exist, then 2.1 follows for all even $n$. It may be easier to first prove or disprove 4.15.
Acknowledgment

### v2

@@ 5 Power functions
If $F(x)=x^{d}$ is APN and $n$ is odd, we can determine the exact exclude multiplicities of $\mathcal{G}_{F}$ contained in $\left\{0\right\}\times\mathbb{F}_{2^{n}}^{\ast}$, but first we describe the sum $\sum_{x\in\mathbb{F}_{2^{n}}}\gamma_{F}(x+a,x^{d}+b)$ in the following way to make our computation straightforward.
@@
In [28, Conjecture 21], Kaleyski conjectured that if $F$ is an APN power function, the two values $\sum_{y\in\mathbb{F}_{2^{n}}}\delta_{F}(y,F(y)+1)$ and $\sum_{y\in\mathbb{F}_{2^{n}}}\delta_{F}(y+1,F(y))$ are equal to $\sum_{y\in\mathbb{F}_{2^{n}}}\delta_{x^{3}}(y,y^{3}+1)$. By Proposition 3.3, this is equivalent to $\operatorname{mult}_{\mathcal{G}_{F}}(0,1)=\operatorname{mult}_{\mathcal{G}_{F}}(1,0)=\operatorname{mult}_{\mathcal{G}_{x^{3}}}(0,1)$. Therefore, Proposition 5.3 proves Kaleyski’s conjecture in the case of $n$ odd, but the case of $n$ even still remains open.
@@ 6 Lower bounds on distances from the APN inverse function
Kloosterman sums can also be used to describe the number of rational points on ordinary elliptic curves over $\mathbb{F}_{2^{n}}$. Recall that an algebraic plane curve $\mathcal{C}$ over $\mathbb{F}_{2^{n}}$ is defined by the equation $p(x,y)=0$ for some irreducible polynomial $p(x,y)$ over $\mathbb{F}_{2^{n}}$. If $\widetilde{p}(x,y,z)$ is the associated homogenized polynomial to $p(x,y)$ we denote the solution set to $\widetilde{p}(x,y,z)=0$ as $\widetilde{C}$. We say that the number of rational points of $C$ is the size of $\widetilde{C}$. Also recall that a point $(x,y,z)\in\widetilde{C}$ is singular if all partial derivatives of $\widetilde{p}(x,y,z)$ vanish at $(x,y,z)$. If there are no singular points, we define the genus of $\mathcal{C}$ as $g=\frac{(d-1)(d-2)}{2}$, where $d$ is the degree of $\widetilde{p}(x,y,z)$. If the genus of $\mathcal{C}$ is $1$, then we say that $\mathcal{C}$ is elliptic. Moreover, we say that $\mathcal{C}$ is supersingular if the coefficient of xyzxyz is zero in $\widetilde{p}(x,y,z)$, otherwise we say that $\mathcal{C}$ is ordinary (cf. [33]).
@@
$\operatorname{mult}_{\mathcal{G}_{F}}(a,b)\geq\left\lceil\frac{2^{n}-5-\max_{a\in\mathbb{F}_{2^{n}}^{\ast}}|K_{n}(a)-1|}{6}\right\rceil.$ (10)
@@
Let $\mathcal{E}=\{(x,y)\in\mathbb{F}_{2^{n}}^{2}:q(x,y)=0\}$, and let $G=\left\{(x,y)\in\mathcal{E}:0\notin\left\{x,y,x+y+1\right\}\right\}$. Note that for all $(x,y)\in\mathbb{F}_{2^{n}}^{2}$ such that $x,y,x+y+1\neq 0$, we have $p(x,y)=0$ if and only if $q(x,y)=0.$ Hence $|D|=|G|$.
@@
Let $\widetilde{\mathcal{E}}=\{(x,y,z):\tilde{q}(x,y,z)=0\}$ denote the projective curve defined by $\tilde{q}$. By considering the partial derivatives of $\tilde{q}$, it is easy to see that $\mathcal{E}$ is smooth, and we know that $\mathcal{E}$ is elliptic because its genus is $g=\frac{(3-1)(3-2)}{2}=1$. Moreover, $\mathcal{E}$ is ordinary because xyzxyz has a nonzero coefficient in $\tilde{q}$. Note that the projective points of $\widetilde{\mathcal{E}}$ include three additional points, corresponding to solutions of the form $(x:y:0)$ that satisfy $\widetilde{q}(x,y,0)=bx^{2}y+bxy^{2}=bxy(x+y)=0$. There are three such cases: $x=0,y=0$ and $x+y=0,$ corresponding to the points $(1:0:0),(0:1:0),(1:1:0)\in\widetilde{\mathcal{E}}$ respectively. So $\widetilde{\mathcal{E}}=\#\mathcal{E}+3$.
@@ 7 On the existence of linear structures of $\gamma_{F}$
In [15], two classes of APN functions were given such that $\gamma_{F}$ does not admit non-trivial linear structures. The first is when $F$ is an APN power function, and the second is when $F$ is such that there exists some $x\in\mathbb{F}_{2}^{n}$ with the property that the size of $\left\{(a,b)\in(\mathbb{F}_{2}^{n})^{2}:D_{a}D_{b}F(x)=w\right\}$ does not depend on $w\in\mathbb{F}_{2}^{n}\setminus\left\{0\right\}$.
We prove that if $F$ is a plateaued APN function with $n$ even, then $\gamma_{F}$ does not admit nontrivial linear structures, forming the third known class of APN functions such that $\gamma_{F}$ has no nontrivial linear structures (excluding AB functions as $\gamma_{F}$ is bent in this case and therefore has balanced nonzero derivatives). Let us rephrase the problem of $\gamma_{F}$ having linear structures in terms of exclude multiplicity when $F$ is a plateaued APN function.
Lemma 7.1.
@@
The following result is then an immediate corollary of the fact that plateaued APN functions cannot have an algebraic degree equal to $n$, see [7].
@@
Assume $n\geq 4$ is even. Let $F\colon\mathbb{F}_{2}^{n}\to\mathbb{F}_{2}^{n}$ be a plateaued APN function. Then $\gamma_{F}$ has no nontrivial linear structures.
@@
Assume that $F$ is plateaued, and by way of contradiction assume that $(0,F(0)+b)$ has an exclude multiplicity of $\frac{2^{n}-1}{3}$ for some $b\neq 0$. Then $\mathbb{F}_{2}^{n}\setminus\{0\}$ decomposes into $\frac{2^{n}-1}{3}$ disjoint triples $\{x,y,z\}$ with
$(x+y+z,F(x)+F(y)+F(z))=(0,b+F(0)).$
Summing over all triples gives $\sum_{x\in\mathbb{F}_{2}^{n}\setminus\left\{0\right\}}(x,F(x))=(0,b+F).$ Moreover, since $F$ has algebraic degree strictly less than $n$, we know $\sum_{x\in\mathbb{F}_{2}^{n}}F(x)=0$, implying $\sum_{x\in\mathbb{F}_{2}^{n}\setminus\left\{0\right\}}(x,F(x))=(0,F(0))$. Therefore, $b=0$, a contradiction. Thus, no exclude point of $\mathcal{G}_{F}$ can have multiplicity equal to $\frac{2^{n}-1}{3}$ by Proposition 4.4, and the result follows from Lemma 7.1. ∎
As shown above, there is an equivalence of $\gamma_{F}$ having a nontrivial linear structure and $\mathcal{G}_{F}$ having an exclude point of multiplicity $\frac{2^{n}-1}{3}$ when $F$ is plateaued APN, and we showed that this is an impossibility. In the case of $F$ also being quadratic APN, we can then demonstrate that the ortho-derivative of $F$ cannot have trivial component functions.
@@
Suppose $n\geq 4$ is even, and let $F\colon\mathbb{F}_{2}^{n}\to\mathbb{F}_{2}^{n}$ be a quadratic APN function. Then $\pi_{F}$ has no components of weight 0, or equivalently, the image of $\pi_{F}$ is not contained in a linear hyperplane.
@@
Apply Lemma 7.1 and Proposition 7.2 to eq. 9. ∎
We now apply the inequality $\operatorname{wt}(b\cdot F)\geq\operatorname{\mathcal{NL}}(\pi_{F})$ for $b\neq 0$ to eq. 9. In particular, we have
$\operatorname{mult}_{\mathcal{G}_{F}}(a,b)\leq\frac{2^{n}-1-\operatorname{\mathcal{NL}}(\pi_{F})}{3}$ (13)
@@
Suppose $n\geq 4$ is even. Let $F\colon\mathbb{F}_{2}^{n}\to\mathbb{F}_{2}^{n}$ be a plateaued APN function with $F(0)=0$. Then for any $b\neq 0$, the weight of $D_{(0,b)}\gamma_{F}$ is divisible by $6\cdot 2^{n}$ and satisfies
@@
By Proposition 4.14, we know $\operatorname{mult}_{\mathcal{G}_{F}}(0,b)=2k+1$ for some non-negative integer $k\geq 0$. Therefore, by Proposition 4.16, we have $6\cdot 2^{n}(2k+1)=2^{n}(2^{n}-2)+2^{2n}-2\operatorname{wt}(D_{(0,b)}\gamma_{F})$. So $\operatorname{wt}(D_{(0,b)}\gamma_{F})=2^{n}(2^{n}-6k-4)$. Since $n$ is even, $2^{n}-4$ is divisible by $6$, and so we have $\operatorname{wt}(D_{(0,b)}\gamma_{F})\equiv 0\mod 6\cdot 2^{n}$.
By Theorem 4.13, we have $\operatorname{mult}_{\mathcal{G}_{F}}(a,b)\geq 2^{\frac{n}{2}-1}-1$ for all $(a,b)\notin\mathcal{G}_{F}$. Moreover, we have $\operatorname{mult}_{\mathcal{G}_{F}}(a,b)\leq\frac{2^{n}-1}{3}$ by [20, Corollary 3.4]. So, for $b\neq 0$, we have $2^{\frac{n}{2}-1}-1\leq\operatorname{mult}_{\mathcal{G}_{F}}(0,b)\leq\frac{2^{n}-1}{3}$, and the result follows. ∎
@@ 8 Open problems
For $n$ even, find an example of a non-plateaued APN function $F\colon\mathbb{F}_{2}^{n}\to\mathbb{F}_{2}^{n}$ whose graph has an exclude point of even multiplicity, or prove such a function does not exist. Note that if such a function does not exist, then 2.1 follows for all even $n$.
Declarations
Ethics approval and consent to participate:
Not applicable.
Consent for publication:
Not applicable.
Availability of data and materials:
This study does not generate or analyze datasets.
Competing interests:
The authors declare they have no competing interests.
Funding:
This research received no external funding.
Authors’ contributions:
All authors contributed equally to this work.
Acknowledgments: